OpenStudy (anonymous):

Which is the simplified form of 18 times q to the fourth power times r all over 6 times q cubed times r to the fifth power.? 3 times r to the fourth power all over q. 12 times r to the fourth power all over. 3 times q all over r to the fourth power. 12 times q all over r to the fourth power.

OpenStudy (anonymous):

18rq^4 /6q^3 *r^5 ?

OpenStudy (anonymous):

Split it up: (18/6) (r/r^5)(q^4/q^3)

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

How to do one of the terms: r/ r^5 = r^1 * r^-5 = r^-4 = 1/r^4

OpenStudy (anonymous):

idk, i dont get this question

OpenStudy (anonymous):

18 - r = 18-r. I can't get a good answer like 4 with that. Since they're multiplied, you can split it up like I did: (18/6) (r/r^5)(q^4/q^3). 18 can be divided by 6, r can be divided by r^5, likewise with q^4/q^3.

OpenStudy (anonymous):

The equation is equivalent to: 18 * r * q^4 ---------------- 6 * r^5 * q^3

OpenStudy (anonymous):

Do you know how to divide exponents?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay, when you're dividing exponents like: x^5/x^3 , what you're really doing is x^(5-3). X is the same for both of them, so you can do that. They won't divide if it's something like x^4/y^3. x^(5-3) = x^2. If it was x^3/x^5, you would go x^(3-5) = x^-. 2. x^-2 = 1/x^2. Can you solve it now?

OpenStudy (anonymous):

Was I confusing in explaining it?

OpenStudy (anonymous):

I can't just give you the answer, you have to be able to learn it yourself.

OpenStudy (anonymous):

yeah... i still dont get it thats why. but its fine ill just teach myself.

OpenStudy (mertsj):

Divide 18 by 6. What did you get?

OpenStudy (anonymous):

3

OpenStudy (mertsj):

Good. Now watch this:

OpenStudy (mertsj):